Hawk Host Status

Dallas, TX

Dallas Datacenter

Time Status Message
2 weeks, 4 days ago Account migrations of the Mercury system to DAL112 are completed at this time.
3 weeks ago Account migrations of the Mercury system to DAL112 are underway. Please contact our support team if you have any questions or issues.
1 month, 1 week ago The failing disk on dal-sd001 has been replaced and i/o latency is once again normal.
1 month, 1 week ago The dal-sd001 system has a disk that is currently failing resulting in abnormal i/o latency. Our team are currently working on replacing the failing disk.
1 month, 2 weeks ago The cougar system is now back online
1 month, 2 weeks ago The cougar system has had a kernel panic. We are currently investigating
1 month, 3 weeks ago The attack against our Dallas location has been filtered.
1 month, 3 weeks ago Our Dallas location is currently the target of a large distributed denial of service attack. We are currently working on filtering the attack.
3 months ago The machine is now online.
3 months ago The Mercury server is currently having issues - we're investigating.
3 months, 2 weeks ago The issue has been resolved and MySQL is back online
3 months, 2 weeks ago The dal008 system is showing MySQL InnoDB corruption. We are working diligently to resolve the problem.
6 months ago The se001 maintenance should now be completed.
6 months ago We are currently performing maintenance on the se001 SpamExperts cluster. The service may be slow or even inaccessible during this time.
6 months, 3 weeks ago The failed drive has been replacing and the new one is rebuilding.
6 months, 3 weeks ago The lightning server has had one of it's drives fail. We are in the process of replacing it now.
7 months, 2 weeks ago At this time mirage is back online. If you have any further issues or questions please contact our support team.
7 months, 2 weeks ago We anticipate the restore completing within 6 hours. At that time we will begin starting system services and restoring access to customer accounts.
7 months, 2 weeks ago The full restore to new hardware is still underway.
7 months, 2 weeks ago We have encountered unexpected issues recovering system data. As a result we are proceeding with a full restore from our backup servers to new hardware. Please contact our support team for additional updates.
7 months, 2 weeks ago Unfortunately the fsck is still running at this time.
7 months, 2 weeks ago Due to large amounts of corruption the fsck is still running.
7 months, 2 weeks ago We are currently running an fsck on mirage to repair any filesystem corruption.
7 months, 2 weeks ago The hardware issue on mirage is still being actively investigated. Thank you for your patience.
7 months, 2 weeks ago We are still investigating the root cause of the hardware issue causing downtime on mirage.
7 months, 2 weeks ago We are currently investigating an issue with the mirage server.
7 months, 3 weeks ago All services should be back online at this time.
7 months, 3 weeks ago We are currently investigating a network connectivity issue.
8 months, 1 week ago The dal007 system is now back online
8 months, 1 week ago The dal007 system is currently non responsive. We are currently investigating
8 months, 4 weeks ago Network connectivity has been restored
8 months, 4 weeks ago The algae system has lost network connectivity. We are currently investigating
9 months, 2 weeks ago The failed drive has been replaced.
9 months, 2 weeks ago The mars system is currently experiencing abnormally high disk latency due to a failed drive. We are currently working on replacing the failed disk.
10 months, 3 weeks ago The failed drive has been replaced.
10 months, 3 weeks ago The dal009 system is currently experiencing abnormally high disk latency due to a failed drive. We are currently working on replacing the failed disk.
10 months, 3 weeks ago The mercury system is now back online.
10 months, 3 weeks ago The mercury system has had one of it's disks fail. We will be taking the system offline to replace it.
11 months ago The failed drive on dal009 has now been replaced
11 months ago The dal009 system is currently experiencing abnormally high disk latency due to a failed drive. We are currently working on replacing the failed disk.
11 months, 2 weeks ago This should now be resolved.
11 months, 2 weeks ago We're currently experiencing a large DDoS attack causing intermittent network issues. We're working on mitigating it at this time.
1 year ago The maintenance from yesterday was completed.
1 year ago The intermittent issues have been resolved. Our networking team continue to perform further maintenance.
1 year ago We're actively investigating intermittent network issues for some systems in our Dallas location
1 year ago We will be performing maintenance on our network and servers in Dallas so there may be intermittent outages. We'll post updates upon completion.
1 year, 1 month ago The failed drive on skyline has now been replaced
1 year, 1 month ago The Skyline system is currently experiencing abnormally high disk latency due to a failed drive. We are currently working on replacing the failed disk.
1 year, 2 months ago The failed drive on skyline has now been replaced
1 year, 2 months ago The Skyline system is currently experiencing abnormally high disk latency due to a failed drive. We are currently working on replacing the failed disk.
1 year, 8 months ago Network traffic has been restored to normal. We'll be monitoring the situation for any latent issues.
1 year, 8 months ago Currently there is some network loss upstream causing some slowness depending on your ISP. This is still being investigated.
1 year, 8 months ago There are reported network issues. At this time we're investigating.
1 year, 8 months ago Network connectivity has been restored. Our networking team will continue to closely monitor the network.
1 year, 8 months ago We've received reports of continued connectivity issues to Dallas. Our networking team are still working on the issue.
1 year, 8 months ago Network connectivity has been restored. Our networking team will continue to closely monitor the network.
1 year, 8 months ago Our networking team are still working on the routing issues in Dallas.
1 year, 8 months ago Our Dallas location is currently experiencing routing issues. Our networking team are investigating.